OpenStudy (vortish):

find the maximum y value on the graph y=f(x) f(x)=-x^2=8x+6

OpenStudy (anonymous):

@vortish f(x) = -x^2 = 8x+6 ??????? is it right?

OpenStudy (vortish):

no it is f(x)=-x^2+8+6

OpenStudy (anonymous):

ok take the first derivative of f(x) and put the derivative equal to zero and tell me what u r getting?

OpenStudy (vortish):

i dont understand that can you draw it and then i will be understand what your asking pls

OpenStudy (anonymous):

cool it is given f(x) = -x^2+8x+6 can u do the differentiation of f(x) w.r.t x ?

OpenStudy (vortish):

no I always had trouble with these expressions

OpenStudy (anonymous):

fine as we know d/dx ( x^n) = n x^(n-1) and d/dx(constant) = zero so f ' (x) = -2x+8 did u follow this step?

OpenStudy (vortish):

yup

OpenStudy (anonymous):

now put f '(x) = o and tell me what is the value of x u r getting?

OpenStudy (vortish):